If a gas of volume 0.06 m3




and at pressure 100kPa is compressed quasistatically according to




pV2 =C until the volume become 0.02m3




, determine the final pressure and the work transfer.

1
Expert's answer
2022-03-10T08:44:03-0500

pVn = C

n = 2

(V2/V1)n = p1/p2

(0.02 m3/0.06 m3)2 = 100 kPa /p2

p2 = 900 kPa

W = p1V1ln(p1/p2) = 100 × 103 Pa × 0.06 m3 × ln(100/900) = 13 183 J = 13.183 kJ